Given g(x)= (x^2)/(x^2-4)

a) Find all critical values for 𝑔(π‘₯).


b) Find 𝑔 β€²β€²(π‘₯). Then, use your answers from part (a) and the 2nd derivative test to determine if each critical value represents a relative maximum, minimum, or neither


c) On what interval(s), if any, is 𝑔(π‘₯) concave down. Justify.


1
Expert's answer
2021-01-29T05:09:41-0500

Consider the function "g(x)= \frac{x^2}{x^2-4}"


a) Let us find the derivative of "g(x):"


"g'(x)=\frac{2x(x^2-4)-x^22x}{(x^2-4)^2}=-8\frac{x}{(x^2-4)^2}"


In the points "-2" and "2" the function "g(x)" doesn’t exist. If "g'(x)=0", then "x=0." Therefore, "x=0" is a critical value for "\ud835\udc54(\ud835\udc65)".


b) Let us find "\ud835\udc54''(\ud835\udc65)":


"g''(x)=(-8\frac{x}{(x^2-4)^2})'=-8\frac{(x^2-4)^2-x\cdot 2(x^2-4)2x}{(x^2-4)^4}=
-8\frac{x^2-4-4x^2}{(x^2-4)^3}=8\frac{3x^2+4}{(x^2-4)^3}"


Since "g''(0)=-0.5<0", we conclude that the critical value "x=0" represents a relative maximum.


c) Since "g''(x)<0" for "x\in(-2,2)", we conclude that on the interval "(-2,2)" the function "\ud835\udc54(\ud835\udc65)" is concave down.